OpenStudy (jayden):

Solve for x: 1 over 3 (2x - 8) = 4 (1 point) 2 6 10

alones (alones):

\(\frac{ 1 }{ 3 }(2x - 8) = 4\)?

OpenStudy (jayden):

yes

alones (alones):

Alright simplifying would be your first step.

alones (alones):

After you should add \[\frac{ 8 }{ 3 }\]

alones (alones):

So than multiply each side by \(\frac{ 3 }{ 2 }\).

alones (alones):

Anyways so what would your answer be?